OpenStudy (anonymous):

Derivative question. let y=e^(9sin(9t)) dy/dt=??

OpenStudy (kropot72):

\[\frac{dy}{dx}=\frac{d\ e ^{9\sin (9t)}}{d\ 9\sin (9t)}\times \frac{d\ 9\sin (9t)}{dx}\]

OpenStudy (anonymous):

what is LXIX in roman numerals HELP !!!!! I REALLY NEED HELP and PLEASE EXPLAiN HOW U GET THE ANSWER